OpenStudy (anonymous):

@imdestroyer69 At the movie theatre, child admission is $6.20 .and adult admission is $9.30. On Tuesday, four times as many adult tickets as child tickets were sold, for a total sales of $954.80. How many child tickets were sold that day?

OpenStudy (anonymous):

9.30*4+6.20=x 954.80/x

OpenStudy (anonymous):

i think then work tat out

OpenStudy (anonymous):

then whatever u get /5 = ur answer

OpenStudy (texaschic101):

6.2c + 9.3a = 954.80 a = 4c 6.2c + 9.3(4c) = 954.8 6.2c + 37.20c = 954.8 43.4c = 954.8 c = 954.8/ 43.4 c = 22 a = 4c a = 4(22) a = 88 check.. 6.2c + 9.3a = 954.80 6.2(22) + 9.3(88) = 954.80 136.40 + 818.40 = 954.80 954.80 = 954.80 (correct) there were 22 children tickets and 88 adult tickets

OpenStudy (anonymous):

i was wrong ig

OpenStudy (anonymous):

ty and its ok